OpenStudy (anonymous):

if you integrate -1/x ... is your answer ln|-x| or is is -ln|x| or is it -ln|-x|??

zepdrix (zepdrix):

\[\large \int\limits -\dfrac{1}{x}dx \quad = \quad -\int\limits \dfrac{1}{x}dx \quad =\quad -\ln|x|+c\] Hmm now you've got me thinking.. whether or not we can bring that negative into the log... Brain... esplode +_+

OpenStudy (anonymous):

thank you. :-)